OpenStudy (anonymous):

totally lost please help, use implicit diffe. to find dy/dx, 2xy+y^2=x+y

OpenStudy (anonymous):

y(x)1-2y/2x+2y-1

OpenStudy (anonymous):

Totally lost sounds bad. Do you know how to differentiate implicitly?.

OpenStudy (anonymous):

Lolz totally

OpenStudy (anonymous):

yes thats the answer but again, im lost, @ no data, no sir =(

OpenStudy (anonymous):

mon is when my profesor went over this and i was snowed in.

OpenStudy (anonymous):

well but maybe you can diferentiante something like this: y = x^2

OpenStudy (anonymous):

using dy/dx?

OpenStudy (anonymous):

yeah!. and what is the do you get by differentiate y = x^2?

OpenStudy (anonymous):

2x/y?

OpenStudy (anonymous):

Mmm that was close look: \[\frac{dy}{dx}=2x\]

OpenStudy (anonymous):

ok i get that.

OpenStudy (anonymous):

It's very important that you know how to get the derivative of explicit functions. By the way, what is the difference between and explict function and an implicit function?

OpenStudy (anonymous):

Ok suppose that we have to get the derivative of yx = x^3

OpenStudy (anonymous):

You have to use the chain rule. Do you know what rule is that?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

f'(x)+f(g(x))?

OpenStudy (anonymous):

Mmm not exactly.

OpenStudy (anonymous):

the chain rule can be written as \[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}\]

OpenStudy (anonymous):

Let's apply this chain rule to the equation a wrote before.

OpenStudy (anonymous):

\[yx=x^3\]

OpenStudy (anonymous):

taking the derivative on both sides of the equation.

OpenStudy (anonymous):

\[\frac{d}{dx}(yx)=\frac{d}{dx}(x^3)\]

OpenStudy (anonymous):

\[\frac{dy}{dx}\frac{dx}{dx}=3x^2\]

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

No, it's an example. You need to apply the same rule to your problem.

OpenStudy (anonymous):

no i mean to ur example, is that the answer?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

kk.ima give a go, thanks for the time, i really do appreciate it

OpenStudy (anonymous):

You're welcome!